OpenStudy (anonymous):

Find the 7th term of a geometric sequence with t1 = 6 and r = 3

OpenStudy (anonymous):

A 652 B 1458 C 486 D 4374

hartnn (hartnn):

the 'n' th term of a GP is given by t1*r^(n-1) where t1 is first term =6 and r=3 so put n=7 in that equation....

OpenStudy (anonymous):

b?

OpenStudy (helder_edwin):

follow what hartnn gave you \[ \large t_7=t_1\cdot3^6= \]

hartnn (hartnn):

so whats 6*3^6 ?

OpenStudy (anonymous):

18^6?

hartnn (hartnn):

ohh,its actually 6*(3^6)=6*729=4347